OpenStudy (anonymous):

Use the equation 4x + 10 = 32 for parts a - d a. Add 4 to each side of the equation. Tell in your own words, how you know that the two sides of the equation are still equal. B. How are the solutions of the original equation and the equation inn part a related? What does the term equivalent equation mean? C. Use subtraction, multiplication, and division to form three equations equivalent to the original equations. D. Solve the original equation. Explain each step in finding the solution? I hope someone can help me I am so confused

OpenStudy (dogzcatz):

Original equation is x=11/2 or x=5.5

OpenStudy (dogzcatz):

steps to original problem. 4x=10=32 4x=32-10 4x=22 x=22\[\div\]4 x=22/4

OpenStudy (anonymous):

Thank you for breaking it down to me.

OpenStudy (anonymous):

I am still a little confused though